# Taming Latent Diffusion Models to See in the Dark

### 10 Dec 2023 v2

@@ Abstract
Enhancing a low-light noisy RAW image into a well-exposed and clean sRGB image is a significant challenge in computational photography. Due to the limitation of large-scale paired data, prior approaches have difficulty in recovering fine details and true colors in extremely low-light regions. Meanwhile, recent advancements in generative diffusion models have shown promising generating capabilities, which inspires this work to explore generative priors from a diffusion model trained on a large-scale open-domain dataset to benefit the low-light image enhancement (LLIE) task. Based on this intention, we propose a novel diffusion-model-based LLIE method, dubbed LDM-SID. LDM-SID aims at inserting a set of proposed taming modules into a frozen pre-trained diffusion model to steer its generating process. Specifically, the taming module fed with low-light information serves to output a pair of affine transformation parameters to modulate the intermediate feature in the diffusion model. Additionally, based on the observation of dedicated generative priors across different portions of the diffusion model, we propose to apply 2D discrete wavelet transforms on the input RAW image, resulting in dividing the LLIE task into two essential parts: low-frequency content generation and high-frequency detail maintenance. This enables us to skillfully tame the diffusion model for optimized structural generation and detail enhancement. Extensive experiments demonstrate the proposed method not only achieves state-of-the-art performance in quantitative evaluations but also shows significant superiority in visual comparisons. These findings highlight the effectiveness of leveraging a pre-trained diffusion model as a generative prior to the LLIE task. Project page: https://csqiangwen.github.io/projects/ldm-sid/.
{strip}
Figure 1: An extremely low-light real-world image from SID Dataset [4] with a ×$\times 300$ amplification ratio. The image processed by our proposed method exhibits clear improvement in structural content, showcasing finer details compared to previous works. The description generated for each image by InstructBLIP [6] also suggests that our result provides higher quality to downstream tasks.
@@ 1 Introduction
Low-light image enhancement (LLIE) is a crucial task in computational photography, particularly for RAW images. Enhancing a low-light noisy RAW image into a well-exposed and clean sRGB image not only requires addressing noise reduction but also grapples with challenges related to color correction, among other image signal processing operations.
In recent years, learning-based methods [4, 14, 38, 32, 30, 37, 5, 12, 43, 11] have emerged as the dominant approach and have made remarkable progress in the LLIE task. Several studies [4, 38, 22, 5, 12, 34, 31, 36] collect datasets with real-world noisy-clean data pairs for training these deep models. Nevertheless, due to the limited range of scenes that these datasets cover, learning-based methods inevitably suffer from noticeable artifacts, e.g., over-smooth and color distortion problems, in extremely dark regions.
An intuitional way to improve the performance of learning-based methods is to increase the number of paired training data. To avoid the challenges associated with laborious real-world data collection, some works [9, 32, 28, 1, 3, 38] focus on designing noise models to synthesize realistic noisy-clean data pairs to train LLIE methods on much more data which improves the generalization. However, the gap between the synthetic and real-world noise distribution still impedes LLIE methods to achieving satisfying restoration and denoising performance in challenging cases as mentioned in [41, 17, 16].
With the rapid development of generative models, especially diffusion models, some works [8, 27, 39] demonstrate the generative prior from a pre-trained diffusion model shows promising performance in image restoration tasks. Instead of burdensomely training a diffusion model from scratch [42, 30, 18, 13] on a specific dataset, a diffusion model trained on a large-scale open-domain dataset still shows powerful generating ability in a target task.
@@
Inspired by recent works [8, 27, 39], we propose a novel LLIE method, LDM-SID, based on a frozen pre-trained diffusion model [20]. To fully exploit the generative prior, instead of directly finetuning the parameters of the diffusion model, LDM-SID aims at inserting a set of trainable taming modules, whose key part is the spatial feature transformation (SFT) layer [29] to modulate intermediate features of the diffusion model for the LLIE task. Meanwhile, with an observation that there are dedicated generative priors in different portions of the diffusion model, we propose to employ 2D discrete wavelet transforms to divide the LLIE task into two parts: low-frequency content generation and high-frequency detail maintenance. To be specific, LDM-SID employs 2D discrete wavelet transforms on the low-light noisy RAW image to extract its low-frequency and high-frequency sub-bands. Then, the low-frequency sub-band is fed into taming modules to steer the structural content generation while the high-frequency is adopted to maintain the precise and clear details in the final sRGB output. Feeding low- and high-frequency subbands separately allows us to skillfully exploit dedicated generate priors in different portions of the diffusion model. By training the proposed taming modules on a small-scale LLIE dataset, LDM-SID finds an optimized balance between image fidelity and perceptual quality, which allows it to achieve state-of-the-art LLIE performance by generating high-quality and artifact-free images.
To sum up, our primary contributions are:
@@
  • We propose a novel diffusion-model-based LLIE method, LDM-SID, which is simple but effective. Instead of training a diffusion model on a specific LLIE dataset, LDM-SID only needs to insert a set of trainable taming modules into a pre-trained diffusion model to generate high-quality outputs. Keeping the pre-trained diffusion model frozen allows LDM-SID to fully exploit generative priors from the diffusion model trained on a large-scale open-domain dataset.
@@
  • To maximally exploit dedicated generative priors in different portions of the diffusion model, we further propose to divide the LLIE task into low-frequency content generation and high-frequency detail maintenance by applying 2D discrete wavelet transforms (DWT) on the input RAW image. This ensures that LDM-SID achieves an optimized balance between generating accurate structural content and achieving fine-grained detail enhancement with noise reduction, resulting in improved image quality.
@@
  • Extensive evaluations conducted on four widely-used real-world datasets show the proposed method not only achieves state-of-the-art performance according to quantitative metrics but also significantly outperforms previous works in perceptual quality. Our LDM-SID excels in challenging regions by producing outputs without any remaining noise, highlighting its superiority over prior works.
Figure 2: The overview of our proposed LLIE method, LDM-SID. The Bayer RAW image is processed by similar operations mentioned in [4]. A series of 2D discrete wavelet transforms (DWT) is applied to the processed image for capturing the low-frequency (LL) and high-frequency subbands (LH, HL, HH). The low-frequency subband (LL) serves to modulate the feature at each layer in the LDM. Specifically, each feature has its corresponding taming module, whose key part is an SFT layer, to map the sub-band into a pair of scale $\gamma$ and shift $\beta$ parameters. Similar to the low-frequency taming, the features of the decoder $\mathcal{D}$ are modulated by another set of taming modules, where the LL sub-band is mapped to the scale $\gamma$ and the concatenation of LH, HL, HH is mapped to the shift $\beta$. All parameters from the pre-trained Stable Diffusion are frozen and only taming modules are trainable.
@@ 2 Related Work
2.1 Low-light Image Enhancement
For enhancing low-light RAW images, numerous studies have showcased the promising performance of deep learning-based methods. Several studies [4, 30, 37, 5, 12, 43, 11] propose supervised learning-based methods that leverage real-world paired data for training. Chen et al[4] propose the first RAW dataset, SID, which contains 5096 RAW data pairs and they designed a U-Net-based method to tackle the LLIE problem. Zhu et al[43] propose an edge-enhance multi-exposure fusion network to improve the performance of a learning-based method in the LLIE task. Wang et al[30] propose the ExposureDiffusion which offers an effective approach for reducing noise in low-light images by leveraging the principles of the diffusion model [10]. Although the supervised learning-based methods have shown promising performance, they still suffer from restoring background with noticeable artifacts in heavily noisy areas due to the limited range of scenes that the dataset covers. Since extending the scale of real-world datasets is quite challenging, diverse noise models [32, 38, 9, 28, 1, 3] are proposed to generate realistic noise for synthesizing more low/normal-light data pairs. Physics-based noise modeling [9, 32, 41] has been commonly used in previous works for noise synthesis in low-light RAW images. To improve the accuracy of the unknown noise synthesis, some works [1, 3, 38] focus on designing learning-based noise modeling. However, the gap between the synthetic and real-world noise distributions impedes learning-based methods to achieve satisfactory performance in real-world photography.
@@ 2.1 Low-light Image Enhancement
In this work, instead of training a method on a large-scale synthetic dataset to improve the restoration performance in the LLIE task, we propose to exploit generative priors from a powerful text-to-image diffusion model [20], which is pre-trained on a large-scale sRGB dataset, LAION-5B [21] containing 5.85 billion text-image pairs. With the help of our proposed taming modules, the proposed method LDM-SID solely trains a set of taming modules on the existing small-scale LLIE dataset (e.g., SID-Sony [4]) to tame the pre-trained diffusion model instead of finetuning any parameters in it.
2.2 Prior for Low-light Image Enhancement
Instead of relying on more data pairs to enhance the restoration performance, several studies [33, 8, 35] have focused on incorporating priors into their methods to improve the LLIE performance. Wu et al[33] propose to incorporate the semantic priors into the LLIE method. This method achieves more natural results compared to previous works by integrating image features with semantic features extracted from a pre-trained segmentation network. Xu et al[35] propose to take an SNR (Signal-to-Noise-Ratio) prior to their method to achieve better LLIE performance. Their method first calculates an SNR map from the input noisy image and its corresponding clean image which is generated by a no-learning-based denoising operation. The SNR map is fed into an SNR-aware transformer [25] to guide feature fusion. Most recently, Fei et al[8] propose to utilize the generative prior from a pre-trained diffusion model [7] for unified image restoration, including the LLIE task. By optimizing a carefully designed degradation model, their method regards the input low-light image as the guidance to control the generating process. Since the generated image is supervised by the input low-light image, the heavy noise from the input still remains in the final output. Although the above methods outperform previous works in the LLIE task, they still suffer from noticeable artifacts or remaining noise. Additionally, sRGB input is required by these methods to exploit corresponding priors, which impedes them from achieving LLIE in RAW space.
In this work, the proposed LDM-SID fully exploits the generative prior from a frozen pre-trained diffusion model [20] without any degradation assumption. The generative prior of the diffusion model trained on a large-scale dataset helps LDM-SID achieve noise-free LLIE performance.
3 Generative Priors in the Diffusion Model
Figure 3: A text-to-image example from Stable Diffusion. The difference map between the latent representation and sRGB output indicates that the latent representation generated by the LDM shows essential structural content while the decoder $\mathcal{D}$ introduces abundant details but scarcely modifies the global structure for the final sRGB output. The latent representation is converted to sRGB visualization by the linear transformation [15].
We utilize the Stable Diffusion [20] as our backbone. Stable Diffusion is a text-to-image generative model trained on the LAION-5B dataset [21] which is approximately 3M times bigger than a widely-used LLIE dataset SID-Sony [4]. It consists of an autoencoder and a latent diffusion model (LDM). Delving into the autoencoder, the encoder $\mathcal{E}$ serves to encode an sRGB image into a latent representation $Z$ while the decoder $\mathcal{D}$ reconstructs the sRGB image from $Z$. The LDM performs the denoising process for the noised latent representation $Z_{t}$, where $t$ denotes the timestep.
Recently, Wang et al[27] propose to exploit generative priors of Stable Diffuiosn to benefit the image super-resolution task. However, they do not delve into the presence of dedicated generative priors within different portions of Stable Diffusion. With the further study of these priors within Stable Diffusion, we have two observations shown in Fig. 3. 1) Given a text prompt, the LDM primarily serves to generate the latent representation which shows high-correlation structural information to the prompt and neglects to include intricate details; 2) During decoding the latent representation to the sRGB output, the decoder $\mathcal{D}$ introduces abundant details to improve the perceptual quality of the output but scarcely modifies the global structure. These facts indicate that the LDM mainly contains the low-frequency generative prior while the decoder $\mathcal{D}$ chiefly has the high-frequency generative prior.
@@ 4 Method
An overview of the proposed method is shown in Fig. 2. Given a low-light noisy RAW image, we first process it to match the illumination to our target sRGB image: $X=(\text{RAW}-\text{Black Level})\times\gamma$, where $\gamma$ represents the user-defined amplification ratio. We aim at taming Stable Diffusion to perform the image signal processing pipeline (ISP), restoration for over-dark areas, and denoising for the processed image.
Based on the above observation that the LDM is adept at generating structural content, which is predominantly dictated by low-frequency information, whereas the decoder demonstrates proficiency in the generation of high-frequency details as shown in Fig. 3, we are inspired to extract different-frequency sub-bands from the input image by 2D discrete wavelet transforms (DWT) and separately feed them into the LDM and decoder $\mathcal{D}$ to exploit dedicated generative priors.

### 20 Mar 2024 v3

# LDM-ISP: Enhancing Neural ISP for Low Light with Latent Diffusion Models

@@ Abstract
Enhancing a low-light noisy RAW image into a well-exposed and clean sRGB image is a significant challenge for modern digital cameras. Prior approaches have difficulties in recovering fine-grained details and true colors of the scene under extremely low-light environments due to near-to-zero SNR. Meanwhile, diffusion models have shown significant progress towards general domain image generation. In this paper, we propose to leverage the pre-trained latent diffusion model to perform the neural ISP for enhancing extremely low-light images. Specifically, to tailor the pre-trained latent diffusion model to operate on the RAW domain, we train a set of lightweight taming modules to inject the RAW information into the diffusion denoising process via modulating the intermediate features of UNet. We further observe different roles of UNet denoising and decoder reconstruction in the latent diffusion model, which inspires us to decompose the low-light image enhancement task into latent-space low-frequency content generation and decoding-phase high-frequency detail maintenance. Through extensive experiments on representative datasets, we demonstrate our simple design not only achieves state-of-the-art performance in quantitative evaluations but also shows significant superiority in visual comparisons over strong baselines, which highlight the effectiveness of powerful generative priors for neural ISP under extremely low-light environments. Project page: https://csqiangwen.github.io/projects/ldm-isp/.
@@ 1 Introduction
Enhancing low-light photos to well-exposed and clean RGB images not only requires reducing a significant amount of noise but also grapples with challenges related to detail and structure recovery, color correction, and careful joint optimization of many other operations. Thus, imaging in extremely low-light environments has been a long-standing challenging problem for modern digital cameras. Standard approaches such as increasing the ISO, extending exposure time, and enlarging the aperture during capturing, all have inevitable limitations and thus cannot solve the problem in practice. Learning-based methods, on the other hand, have been making surprising progress toward effective low-light imaging systems [2, 12, 38, 30, 28, 35, 1, 10, 41, 8, 11]. However, most of the existing approaches learn the low-light enhancement model from relatively limited paired data and the enhanced RGB images often suffer from detail deficiency and color deviation issues. In this paper, we leverage the powerful generative priors from a pre-trained latent diffusion model [21] to enhance the neural ISP for low-light imaging.
Due to the significant practical value of low-light imaging, there have been many approaches to solve the task, which can be categorized into RGB-based enhancement [16, 15, 19, 37, 29, 33, 32] and RAW-based enhancement [2, 1, 30, 11, 28, 12, 38]. RGB data, however, contains much less signal than its unprocessed RAW counterparts, and thus the performance of RGB-based enhancement is limited, especially under extremely low-light conditions. To alleviate this issue, Chen et al. [2] propose to learn the entire ISP pipeline with paired low-light RAW and well-exposed sRGB data. Dong et al. [4] present to simulate the monochrome raw image from the colored raw image to compensate the photon number of the dark raw images. DNF [11] improves the neural ISP through decoupling the noise-to-clean and RAW-to-sRGB process, and utilizing the feature-level information of noise-to-clean as feedback to the RAW-to-sRGB process. Although the learned RAW-based low-light enhancement has been made effective progress, existing approaches still learn the enhancement process on limited pairs of RAW-sRGB data, and thus the resultant models can hardly perform well in extremely low-light cases.
Figure 1: An extremely low-light real-world image from SID Dataset [2] with a ×300 amplification ratio. The image processed by our proposed method exhibits clear improvement in structural content, showcasing finer details compared to previous works.($\dagger$ denotes the RAW-to-sRGB method and $\ast$ denotes the RAW-to-RAW method)
In this work, we propose the LDM-ISP, which leverages the powerful generative priors of the pre-trained latent diffusion model (i.e., Stable Diffusion) to enhance the neural ISP for low light. We use one end-to-end network to simulate the entire ISP pipeline, which learns denoising, color transformations, demosaicing, and detail enhancement all in one network. Designing such an end-to-end ISP is not a trivial problem. First, Stable Diffusion is designed for content generation in the RGB domain and it is questionable how to tailor the pre-trained Stable Diffusion to the RAW domain. A naive way is to fine-tune the entire diffusion model parameters, which could, however, lead to catastrophic forgetting of the abundant priors [18, 6]. Second, the low-light image enhancement task requires different levels of content enhancements, such as structural recovery, detail enhancement, and color correction. It is also challenging to seamlessly merge the ISP enhancement to the different stages of the diffusion denoising process.
@@
To overcome these challenges, instead of directly finetuning the parameters of the latent diffusion model, LDM-ISP aims at inserting a set of trainable taming modules, whose key part is the spatial feature transformation (SFT) layer [27] to modulate the intermediate features of the latent diffusion model. Meanwhile, we observe that there are dedicated generative priors in different portions of the latent diffusion model, and thus we propose to employ 2D discrete wavelet transforms to divide the LLIE task into two sub-tasks: low-frequency content generation and high-frequency detail maintenance. To be specific, LDM-ISP employs 2D discrete wavelet transforms on the low-light noisy RAW image to extract its low-frequency and high-frequency sub-bands. Then, the low-frequency sub-band is fed into taming modules to steer the structural content generation while the high-frequency is adopted to maintain the precise and clear details in the final sRGB output. Feeding low- and high-frequency subbands separately allows us to skillfully exploit dedicated generate priors in different portions of the latent diffusion model. By training the proposed taming modules on a small-scale LLIE dataset, LDM-ISP finds an optimized balance between image fidelity and perceptual quality, which allows it to achieve state-of-the-art LLIE performance by generating high-quality and artifact-free images.
Our contributions are summarized as follows.
  • We are the first to propose an enhanced neural ISP for low light with the powerful generative priors of the latent diffusion model.
@@
  • Instead of training or finetuning a diffusion model on a specific LLIE dataset, the proposed method LDM-ISP only needs to insert a set of trainable taming modules into a pre-trained latent diffusion model to generate high-quality outputs. This approach effectively tailors the latent diffusion model to the RAW domain.
@@
  • To maximally exploit dedicated generative priors in different portions of the latent diffusion model, we further propose to divide the LLIE task into low-frequency content generation and high-frequency detail maintenance through the 2D discrete wavelet transforms (DWT) on the input RAW image.
@@
  • Extensive evaluations conducted on three widely-used real-world datasets show the proposed method not only achieves state-of-the-art performance according to quantitative metrics but also significantly outperforms previous works in perceptual quality.
@@ 2 Related Work
2.1 Learning-based Image Signal Processing
Deep learning has shown its advantages in learning the entire ISP process [23, 17, 36, 40, 31]. DeepISP [23] proposes one of the first CNN-based ISPs to replace the traditional human-tuned ISP operations, including denoising, demosaicking, color correction, white balance, and photo finishing steps. Instead of utilizing one neural network to simulate the entire ISP pipeline, CameraNet [17] presents a two-stage model for better RGB rendering performance. CycleISP [36] learns the bidirectional mapping of RGB-to-RAW and RAW-to-RGB to recover the RAW images from sRGB input. Recently, Xing et al. [31] proposes to utilize one single invertible neural network to learn an invertible ISP that can do both the RGB rendering and RAW reconstruction. However, most of the existing learning-based ISPs are designed for general photo-capturing conditions, and their performance under extremely low-light environments is unsatisfying. In this work, we aim to enhance the neural ISP for low light, with the abundant generative priors of pre-trained Stable Diffusion.
2.2 RAW-based Low-light Image Enhancement
For enhancing low-light RAW images, numerous studies have showcased the promising performance of deep learning-based methods. These previous studies can be divided into two categories: RAW-to-RAW methods [30, 38, 28, 12] and RAW-to-sRGB methods [2, 1, 8, 4, 41, 14, 20]. RAW-to-RAW works focus on proposing denoising strategies on the RAW domain. With carefully designed noise modeling strategies [30, 38] or denoising methods [28, 12], such works provide denoised RAW images and apply a known ISP on them for the final visualization. RAW-to-sRGB works focus on tackling denoising and ISP tasks for low-light images. Recently, such works have shown significantly superior LLIE performance by designing multi-stage methods [11, 4] that tackle noise-to-clean and RAW-to-sRGB tasks in separate stages, compared to single-stage method [2, 1, 20]. This further indicates that there is a domain gap between denoising and ISP tasks, which makes the RAW-to-sRGB task more challenging than the RAW-to-RAW.
Compared to RAW-to-RAW methods, RAW-to-sRGB methods show a distinct advantage when the camera ISP is unknown or confidential. Chen et al[2] propose the first RAW-to-sRGB UNet-based method SID with a RAW dataset, which contains 5096 RAW data pairs for tackling the LLIE problem. However, this method exhibits severe color bias and a noticeable loss of details in extremely dark regions, primarily attributed to its simplistic single-stage network architecture. Zhu et al[41] propose a multi-stage method by fusing multi-exposure images to avoid the color bias issue. Xu et al[32] propose a multi-stage method based on a frequency-based decomposition to improve the detail enhancement. However, both methods suffer from error accumulation across stages [11]. Dong et al[4] propose a multi-stage method MCR that first predicts a monochrome raw image for information restoration and then performs LLIE. To address the challenge of the domain ambiguity issue in simultaneous denoising and ISP, Jin et al[11] propose a method DNF to tackle denoising and ISP tasks in separate stages. Although these multi-stage methods show promising LLIE performance, they still grapple with color correction and denoising in extremely challenging scenarios. The main reason for this shortcoming is that current training datasets [2, 30, 38, 4] contain limited scenes, which impedes the generalization ability of these learning-based methods.
@@ 2.2 RAW-based Low-light Image Enhancement
In this work, instead of collecting a new dataset to improve the LLIE performance in extremely challenging scenarios, we propose to exploit generative priors from a powerful text-to-image latent diffusion model [21], which is pre-trained on a large-scale sRGB dataset, LAION-5B [22] containing 5.85 billion text-image pairs. With the help of generative priors learned from a large-scale open-domain dataset, the proposed method LDM-ISP not only presents satisfying denoising and ISP performance for low-light images, but also exhibits significant superiority in detail maintenance on extremely dark and noisy regions.
2.3 Diffusion Priors for Low-light Image Enhancement
With the rapid progress of diffusion models, the generative priors of diffusion models show promising potential for the LLIE task. Jiang et al[9] and Yi et al[34] propose diffusion model-based methods to perform sRGB-to-sRGB LLIE. Training a diffusion model from scratch on a small-scale LLIE dataset not only hampers the generalization ability of the method but also suffers from heavy computational costs. Meanwhile, Wang et al[28] propose to train a diffusion model from scratch for the RAW-to-RAW LLIE task. Based on the physical noise model, their diffusion model achieves low-light denoising in the RAW domain. Nevertheless, the challenges of the unknown ISP and unsatisfactory generalization problems persist in this method. Fei et al[5] propose to utilize the generative prior from a pre-trained diffusion model [3] for unified image restoration, including the LLIE task. By carefully designing the degradation model, their method regards the input low-light image as the guidance to control the generating process. Since the generated image is supervised by the input low-light image, the heavy noise from the input still remains in the final output. Although the above methods have promising potential for the LLIE task, they still suffer from noticeable artifacts or remaining noise.
In this work, instead of training or finetuning a diffusion model on an LLIE dataset, we insert a set of trainable taming modules into a frozen pre-trained latent diffusion model [21]. The generative prior of the latent diffusion model trained on a large-scale dataset helps the proposed method LDM-ISP achieve remarkable ISP enhancement performance for LLIE in diverse real-world challenging scenarios.
Figure 2: The overview of our proposed LLIE method, LDM-ISP. The Bayer RAW image is processed by similar operations mentioned in [2]. A series of 2D discrete wavelet transforms (DWT) is applied to the processed image for capturing the low-frequency (LL) and high-frequency subbands (LH, HL, HH). The low-frequency subband (LL) serves to modulate the feature at each layer in the UNet. Specifically, each feature has its corresponding taming module, whose key part is an SFT layer, to map the sub-band into a pair of scale $\gamma$ and shift $\beta$ parameters. Similar to the low-frequency taming, the features of the decoder $\mathcal{D}$ are modulated by another set of taming modules, where the LL sub-band is mapped to the scale $\gamma$ and the concatenation of LH, HL, HH is mapped to the shift $\beta$. All parameters from the pre-trained Stable Diffusion are frozen and only taming modules are trainable.
@@ 3 Method
An overview of the proposed method is shown in Fig. 2. Given a low-light noisy RAW image, we first process it to match the illumination to our target sRGB image: $X=(\text{RAW}-\text{Black Level})\times\gamma$, where $\gamma$ represents the user-defined amplification ratio. We aim at taming the latent diffusion model to perform the image signal processing pipeline (ISP) with restoration for over-dark areas and denoising for the processed image.
3.1 Generative Priors in the Latent Diffusion Model
Figure 3: A text-to-image example from a well-known latent diffusion model, Stable Diffusion. The difference map between the latent representation and sRGB output indicates that the latent representation generated by the UNet shows essential structural content while the decoder $\mathcal{D}$ introduces abundant details but scarcely modifies the global structure for the final sRGB output. The latent representation is converted to sRGB visualization by the linear transformation [13].
We utilize the Stable Diffusion [21] as our backbone. Stable Diffusion is a text-to-image latent diffusion model trained on the LAION-5B dataset [22] which is approximately 3M times bigger than the widely-used LLIE dataset SID-Sony [2]. It consists of an autoencoder and a UNet. Delving into the autoencoder, the encoder $\mathcal{E}$ serves to encode an sRGB image into a latent representation $Z$ while the decoder $\mathcal{D}$ reconstructs the sRGB image from $Z$. The UNet performs the denoising process for the noised latent representation $Z_{t}$, where $t$ denotes the timestep.
Recently, a diffusion model-based method [26] is proposed to exploit generative priors of Stable Diffuiosn to benefit the image super-resolution task. However, they do not delve into the presence of dedicated generative priors within different portions of Stable Diffusion. With the further study of these priors within Stable Diffusion, we have two observations shown in Fig. 3. 1) Given a text prompt, the UNet primarily serves to generate the latent representation which shows high-correlation structural information to the prompt and neglects to include intricate details; 2) During decoding the latent representation to the sRGB output, the decoder $\mathcal{D}$ introduces abundant details to improve the perceptual quality of the output but scarcely modifies the global structure. These facts indicate that the UNet mainly contains the low-frequency generative prior while the decoder $\mathcal{D}$ chiefly has the high-frequency generative prior.
Based on the above observation that the UNet is adept at generating structural content, which is predominantly dedicated by low-frequency information, whereas the decoder demonstrates proficiency in the generation of high-frequency details as shown in Fig. 3, we are inspired to extract different-frequency sub-bands from the input image by 2D discrete wavelet transforms (DWT) and separately feed them into the UNet and decoder $\mathcal{D}$ to exploit dedicated generative priors.